# references.bib

### v1

@@ I Introduction
The evolution of multiple input multiple output (MIMO) technologies toward distributed implementations, in a cell-free setting, will serve as a infrastructure to support distributed integrated sensing and communication (ISAC) systems and networked sensing implementations [behdad-multi-static-2023]. While a significant progress has been done in the literature to propose the best strategies for conciliating communication and sensing requirements, a primary concern in the design of the physical layer of ISAC is security and privacy. Particularly, the ability of malicious parties to use sensing information to track and locate targets can lead to the emerging of unprecedented attacks.
While eavesdropping attacks have been addressed in a number of physical layer security-based solutions [9737364], privacy attacks at the physical layer and countermeasures remain barely explored. A privacy attack model is proposed in [da-silva-multi-static-2023], which consider an internal attacker that leverages the signals received from ISAC access points (APs) to estimate the position of targets by inferring the directions the APs originally beamforming their sensing signals toward these targets. Moreover, the work in [10587082] proposes a secure sensing scheme to prevent anauthorized sensing from a malicious communication user via beamforming design.
Therefore, inspired by the attack model in [da-silva-multi-static-2023], this paper proposes an iterative framework of two blocks, precoder design and access point selection. The precoder design aims at maximizing the signal-to-interference-plus-noise ratio (SINR) at the sensing receivers given communication constraints. The access point selection aims at minimizing the mutual information (MI) between the received signal at users and the sensing signal, by rearranging the access points that transmit ISAC signals and the sensing receivers.
@@ II System Model
Consider a centralized cell-free massive MIMO system of $N_{\text{Tx}}$ transmitting ISAC APs, with $M$ isotropic antennas, jointly serving $N_{\text{UE}}$ $K$-antenna users. However, while the user terminals may consist of $K$ antennas, a decoder designed by a centralized unit would attempt to extract the data only meant for the user which is not desirable by the adversarial model described later. Therefore, a decoder design can be disregarded and the multiple antenna users can be modelled as single antenna users receiving additional independent observations of the signal. Additionally, $N_{\text{Rx}}$ $M$-antenna receiver APs are performing sensing of a static target present in the system, which is illuminated by the ISAC transmitters. This system is depicted in 1.
@@
$\begin{split}y_{i}[n]=\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,i}s_{i}[n]}_{\text{Desired Signal}}+\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,k}s_{k}[n]}_{\text{Communication Interference}}\\ +\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,t}s_{s}[n]}_{\text{Sensing Interference}}+\underbrace{n_{i}[n]}_{\text{Noise}},\end{split}$ (1)
@@
$\begin{split}\mathbf{y}_{r}[n]=\sum\limits_{j=1}^{N_{\text{Tx}}}\alpha_{j,r}\sqrt{\beta_{j,r}}\mathbf{a}(\theta_{r})\mathbf{a}(\theta_{j})^{T}\mathbf{x}_{j}[n]\\ +\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{H}_{j,r}\mathbf{x}_{j}[n]+\mathbf{n}[n].\end{split}$ (3)
@@ III Adversary Model
The adversary is assumed to be one of the communication users that maliciously estimates the position of the target by inferring the directions of the beampatterns from the transmitting APs. Different from [da-silva-multi-static-2023], herein we propose a gradient descent method to minimize the mean square error of the distance from a line projected from each known AP location, in the direction of the estimated beampattern. Then, if the target is within a predetermined radius, $R_{\text{GD}}$, from the estimated location, the estimate is deemed successful in determining the target position. Mathematically, the process can be described as
@@
$\text{MSE}(\mathbf{q})=\dfrac{1}{N_{l}}\sum\limits_{i=1}^{N_{l}}||\mathbf{q}-\mathbf{y}_{i}||^{2}$ (5)
@@
Since the peak of each AP’s beampattern is its sensing signal angle, the adversary tries to replicate this beampattern. This can be done by taking advantage of the precoder being a hidden variable within the transmitted signal $\mathbf{x}_{j}$, as seen in (2), and can be studied in [10587082]. Therefore, the outer product of the transmitted signals, i.e, $\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}$, will contain information about the precoder, which in turns allows the adversary to sweep $\mathbf{a}(\theta_{s})^{H}\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\mathbf{a}(\theta_{s})$ over an angular search area, $\theta_{s}$, to estimate the beampattern and finding its peak. Since the locations of the APs are known by the adversary, $\theta_{s}$ is therefore also assumed to be, at least, within the same half-plane as the correct angle.
For estimating the transmitted signal, the expectation maximization (EM) algorithm [moon-expectation-maximization-1996] is used. In this case, the log-likelihood function used in the EM procedure is $\log(p(\widetilde{y}_{a,j};\mathbf{x}_{j}))$, where $p(\widetilde{y}_{a,j};\mathbf{x}_{j})$ is the probability density function of the received signal $\widetilde{y}_{a,j}$ at the adversary $a$, from AP $j$, given the transmitted signal $\mathbf{x}_{j}$.
The EM approach is employed due to considering the actual channel between the adversary and AP $j$, as a latent variable within the likelihood function. For convenience, the channel estimate used in the initial iteration of the EM algorithm is modelled to be the actual channel, with added Gaussian noise to mimic some channel estimation error. Thus, the channel estimate $\widehat{\mathbf{H}}_{a,j}$ is a Gaussian random variable with mean $\mathbf{H}_{a,j}$ and variance $\sigma^{2}_{H}\mathbf{I}$, where $\sigma^{2}_{H}$ is the channel estimation error variance.
Following [da-silva-multi-static-2023] and [zhang-deep-2022], $u$ is introduced as an additional latent variable, allowing the definitions $\mathbf{h}_{a,j}=\{\mathbf{H}_{a,j},u\}$, as well as $\boldsymbol{\theta}=\{\mathbf{x}_{j},\nu\}$ to be made, with $\nu$ being the degree of freedom of an underlying gamma distribution (details are in [zhang-deep-2022]). Thereafter, by establishing $q(\mathbf{h}_{a,j})$ as a variational arbitrary distribution, the maximizable expression can be defined as
@@ III-1 Expectation Step
The expectation step maximizes $\mathcal{F}(\widetilde{y}_{a,j};q,\boldsymbol{\theta})$, with respect to the latent variable, which in this case is $q(\mathbf{h}_{a,j})$. Further, through studying [da-silva-multi-static-2023], $\mathcal{F}(\widetilde{y}_{a,j};q,\boldsymbol{\theta})$ is maximized in this step by setting
$\mathbb{E}_{q}[\mathbf{H}_{a,j}]=\left[\boldsymbol{\Omega}_{\mathbf{H}}\left(\frac{1}{\sigma_{H}^{2}}\widehat{\mathbf{H}}_{a,j}^{T}+\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\frac{\mathbb{E}_{q}[u]}{\sigma_{n}^{2}}\right)\right]^{H},$ (7)
$\boldsymbol{\Omega}_{\mathbf{H}}=\left(\frac{1}{\sigma_{H}^{2}}\mathbf{I}+\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\frac{\mathbb{E}_{q}[u]}{\sigma^{2}_{n}}\right)^{-1},$ (8)
$\mathbb{E}_{q}[u]=\mathbb{E}\left[\mathcal{G}\left(\frac{v+2M_{r}}{2},\frac{v+C}{2}\right)\right],$ (9)
where $M_{r}$ and $M$ is the number of antennas at the receiver and transmitter, respectively. Also, $\mathbb{E}_{q}$ denotes the expected value of the variational distribution $q$ of the variables. Finally, $C$ in (9) is defined as
[…]
It can be noted that there is a dependency between $\mathbb{E}_{q}[u]$ and $\mathbb{E}_{q}[\mathbf{H}_{a,j}]$, which is mitigated by using the previous iteration’s $\mathbb{E}_{q}[u]$ in the calculations, and in the case of it being the first iteration, it is randomly initiated.
@@ III-2 Maximization Step
This step involves determining the estimate of $\mathbf{x}_{j}$ by maximizing $\mathcal{F}(\widetilde{y}_{a,j};q,\boldsymbol{\theta})$ with regards to $\boldsymbol{\theta}$, and it is shown in [zhang-deep-2022] that the transmitted signal from AP $j$, may be determined using
$\begin{split}\widehat{\mathbf{x}}_{j}=\operatorname*{argmin}_{\textbf{x}_{j}}\left[(\widetilde{y}_{a,j}-\mathbb{E}_{q}[\mathbf{H}_{a,j}]\textbf{x}_{j})^{\,H}(\widetilde{y}_{a,j}-\mathbb{E}_{q}[\mathbf{H}_{a,j}]\textbf{x}_{j})\right.\\ +\left.\textbf{x}_{j}^{\,H}\boldsymbol{\Omega}_{\mathbf{H}}M\textbf{x}_{j}\right].\end{split}$ (11)
@@ IV Privacy-Preserving Framework
For a cell-free MIMO ISAC system, we propose a joint, two block, precoder and AP selection framework for privacy-preservation, which utilizes alternating optimization of the two blocks. For a given receiver/transmitter AP configuration, the first block finds the precoding vectors that maximize the SINR at the sensing receivers. The second block minimizes the probability of detection of the target from the adversary, by altering the AP configuration. Henceforth, this probability is denoted $P_{D}$. As a convergence criteria, the procedure stops when two iterations leave the AP configuration unchanged. The first iteration assumes all APs as transmitters.
@@ IV-A Precoder Design
$P_{j,a}\leq P_{max},\forall j,\forall a,$ (12c)
where $\mathbf{W}$ is a $MN_{\text{Tx}}\times(N_{\text{UE}}+1)$ matrix consisting of each transmitter APs $M\times(N_{\text{UE}}+1)$ precoder matrix, denoted $\mathbf{W}_{j}$ for AP $j=1,2,...,N_{\text{Tx}}$. In turn, $\mathbf{W}_{j}$ consists of AP $j$’s concatenated precoder vectors for each user plus the sensing precoder, that is $\mathbf{W}_{j}=\left[\mathbf{w}_{j,1},\mathbf{w}_{j,2},...,\mathbf{w}_{j,N_{\text{UE}}},\mathbf{w}_{j,t}\right]$. Hence, the aggregate precoder matrix $\mathbf{W}=\left[\mathbf{W}_{1}^{T},\mathbf{W}_{2}^{T},...,\mathbf{W}_{N_{\text{Tx}}}^{T}\right]^{T}$ columns represent the total precoding vector for a user from every transmitter AP, where the last column represents the target. Further, $\gamma_{\text{UE}_{i}}$ and $\gamma$ represents the total received SINR per user and the minimum SINR threshold for a user, respectively. $P_{j,a}$ is the transmit power from AP $j$’s antenna $a$, and $P_{max}$ is the maximum allowed transmission power for each antenna. Therefore, $P_{j,a}$ can be computed by taking the norm square of each row of $\mathbf{W}$.
@@
In (13), $N$ denotes the number of symbols that are being transmitted to each user, and, consequently, the number of samples in the sensing probing signal. Moreover, $\mathbf{s}[n]$ is the vector of transmitted symbols for all users and the sensing signal, i.e $\left[s_{1}[n],...,s_{N_{\text{UE}}}[n],s_{s}[n]\right]$.
Since the aim is to optimize $\gamma_{s}$ with respect to $\mathbf{W}$, the Concave-Convex Procedure (CCP) is considered and thus, the objective function must be concave. For this purpose the first order Taylor approximation is applied to the term $\mathbf{W}^{H}\mathbf{A}_{r}\mathbf{W}$ around the result of the previous iteration of the procedure, resulting in
@@
A similar approach is used to transform the communication SINR constraint into a convex function where the function $\gamma_{\text{UE}_{i}}$ is expanded to
@@
$\mathbf{w}_{i}=\left[\mathbf{w}^{T}_{1,i},\mathbf{w}^{T}_{2,i},...,\mathbf{w}^{T}_{N_{\text{Tx}},i}\right]^{T}$.
Again the first order Taylor approximation is applied around the previous iteration’s value of $\mathbf{w}_{i}$. Also, the slack variables $\tau_{d}$ and $\tau_{n}$ for the denominator and numerator in (16), respectively, are introduced. The final expression for the constraint in (12b) is therefore

### v2

@@ I Introduction
The evolution of multiple input multiple output (MIMO) technologies toward distributed implementations, in a cell-free setting, will serve as a infrastructure to support distributed integrated sensing and communication (ISAC) systems and networked sensing implementations [behdad-multi-static-2023]. While significant progress has been done in the literature to propose the best strategies for conciliating communication and sensing requirements, a primary concern in the design of the physical layer of ISAC is security and privacy [qu_privacy_2024]. Particularly, the ability of malicious parties to use sensing information to track and locate targets can lead to the emerging of unprecedented attacks [wang_anti-malicious_2024]. However, methods for combating the security and privacy concerns inherent in ISAC has been proposed. For example [su_secure_2023, 10587082], where the sensing functionality is utilized to locate attackers, and hamper eavesdroppers by artificial noise transmission. While eavesdropping attacks have been addressed in a number of physical layer security-based solutions [9737364], privacy attacks at the physical layer and countermeasures remain barely explored. A privacy attack model is proposed in [da-silva-multi-static-2023], which consider an internal attacker that leverages the signals received from ISAC access points (APs) to estimate the position of targets by inferring the directions the APs originally beamforming their sensing signals toward these targets. Moreover, the work in [10587082] proposes a secure sensing scheme to prevent unauthorized sensing from a malicious communication user via beamforming design.
Note that, different from [da-silva-multi-static-2023], we propose a different version of the attack model. Additionally, we propose an iterative framework of two blocks, precoder design and access point selection, to reduce the impact of the attack. The precoder design aims at maximizing the signal-to-interference-plus-noise ratio (SINR) at the sensing receivers given communication constraints. The access point selection aims at minimizing the mutual information (MI) between the received signal at users and the sensing signal, by rearranging the access points that transmit ISAC signals and the sensing receivers.
@@ II System Model
Consider a centralized cell-free massive MIMO system of $N_{\text{Tx}}$ transmitting ISAC APs, with $M$ isotropic antennas, jointly serving $N_{\text{UE}}$ $K$-antenna users. Decoders designed by the centralized control unit would aim to extract data meant for the user, hence, a decoder design can be disregarded and multiple antennas at the users can be modeled as receiving additional independent observations of the signal. Additionally, $N_{\text{Rx}}$ $M$-antenna receiver APs are performing sensing of a static target present in the system, which is illuminated by the ISAC transmitters. This system is depicted in Figure 1.
@@
$y_{i}[n]=\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,i}s_{i}[n]}_{\text{Desired Signal}}+\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,k}s_{k}[n]}_{\text{Communication Interference}}\\ +\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,t}s_{s}[n]}_{\text{Sensing Interference}}+\underbrace{n_{i}[n]}_{\text{Noise}},$ (1)
@@
$\mathbf{y}_{r}[n]=\sum\limits_{j=1}^{N_{\text{Tx}}}\alpha_{j,r}\sqrt{\beta_{j,r}}\mathbf{a}(\theta_{r})\mathbf{a}(\theta_{j})^{T}\mathbf{x}_{j}[n]\\ +\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{H}_{j,r}\mathbf{x}_{j}[n]+\mathbf{n}[n].$ (3)
@@ III Adversary Model
The adversary is assumed to be one of the communication users that maliciously estimates the position of the target by inferring the directions of the beampatterns from the transmitting APs. Henceforth, for the adversary model, the sample index $n$ is dropped and it is implied that the following methods are repeated for every sample.
Different from [da-silva-multi-static-2023], herein we propose a gradient descent method to minimize the mean square error of the distance from a line projected from each known AP location, in the direction of the estimated beampattern. Then, if the target is within a predetermined radius, $R_{\text{GD}}$, from the estimated location, the estimate is deemed successful in determining the target position. Mathematically, the process can be described as
@@
$\text{MSE}(\mathbf{q})=\dfrac{1}{N_{l}}\sum\limits_{i=1}^{N_{l}}||\mathbf{q}-\mathbf{y}_{i}||^{2},$ (5)
@@
Since the peak of each AP’s beampattern is its sensing signal angle, the adversary tries to replicate this beampattern. This can be done by taking advantage of the precoder being a hidden variable within the transmitted signal $\mathbf{x}_{j}$, as seen in (2), and can be studied in [10587082]. Therefore, the outer product of the transmitted signals, i.e, $\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}$, will contain information about the precoder, which in turns allows the adversary to sweep $\mathbf{a}(\theta_{s})^{H}\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\mathbf{a}(\theta_{s})$ over an angular search area, $\theta_{s}$, to estimate the beampattern and finding its peak [liu_joint_2020]. Since the locations of the APs are known by the adversary, $\theta_{s}$ is therefore also assumed to be, at least, within the same half-plane as the correct angle.
For estimating the transmitted signal, the expectation maximization (EM) algorithm [moon-expectation-maximization-1996] is used. In this case, the log-likelihood function used in the EM procedure is $\log(p(y_{j,a};\mathbf{x}_{j}))$, where $p(y_{j,a};\mathbf{x}_{j})$ is the probability density function of the received signal $y_{j,a}$ at the adversary $a$, from AP $j$, given the transmitted signal $\mathbf{x}_{j}$.
The EM approach is employed due to considering the actual channel, $\mathbf{h}_{j,a}\in\mathbb{C}^{M}$, between the adversary and AP $j$, as a latent variable within the likelihood function. For convenience, the channel estimate used in the initial iteration of the EM algorithm is modelled to be the actual channel, with added Gaussian noise to mimic some channel estimation error. Thus, the channel estimate $\widehat{\mathbf{h}}_{j,a}$ is a Gaussian random variable with mean $\mathbf{h}_{j,a}$ and covariance matrix $\sigma^{2}_{h}\mathbf{I}_{M}$, where $\sigma^{2}_{h}$ is the channel estimation error variance, and $\mathbf{I}_{M}$ is an identity matrix of size $M\times M$.
Following [da-silva-multi-static-2023] and [zhang-deep-2022], $u$ is introduced as an additional latent variable, allowing the definitions $\mathbf{z}_{j,a}=\{\mathbf{h}_{j,a},u\}$, as well as $\boldsymbol{\theta}=\{\mathbf{x}_{j},\nu\}$ to be made, with $\nu$ being the degree of freedom of an underlying Student’s $t$-distribution (details are in [zhang-deep-2022]). Thereafter, by establishing $q(\mathbf{z}_{j,a})$ as a variational arbitrary distribution, the maximizable expression can be defined as
@@ III-1 Expectation Step
The expectation step maximizes $\mathcal{F}(y_{j,a};q,\boldsymbol{\theta})$, with respect to the latent variable, which in this case is $q(\mathbf{z}_{j,a})$. Further, through studying [da-silva-multi-static-2023], $\mathcal{F}(y_{j,a};q,\boldsymbol{\theta})$ is maximized in this step by setting
$\mathbb{E}_{q}[\mathbf{h}_{j,a}]=\left[\boldsymbol{\Omega}_{\mathbf{h}}\left(\frac{1}{\sigma_{h}^{2}}\widehat{\mathbf{h}}_{j,a}+\mathbf{x}_{j}y_{j,a}\frac{\mathbb{E}_{q}[u]}{\sigma_{n}^{2}}\right)\right],$ (7)
$\boldsymbol{\Omega}_{\mathbf{h}}=\left(\frac{1}{\sigma_{h}^{2}}\mathbf{I}_{M}+\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\frac{\mathbb{E}_{q}[u]}{\sigma^{2}_{n}}\right)^{-1},$ (8)
$\mathbb{E}_{q}[u]=\mathbb{E}\left[\mathcal{G}\left(\frac{\nu+2K}{2},\frac{\nu+C}{2}\right)\right],$ (9)
where $\mathbb{E}_{q}$ denotes the expected value of the variational distribution $q$ of the variables, and $\mathbf{I}_{M}$ is an $M\times M$ identity matrix. Finally, $C$ in (9) is defined as
$C=\frac{1}{\sigma^{2}_{n}}\left[\left(y_{j,a}-\mathbb{E}_{q}[\mathbf{h}_{j,a}]^{H}\mathbf{x}_{j}\right)^{\!2}\,+M\mathbf{x}_{j}^{H}\boldsymbol{\Omega}_{\mathbf{h}}\mathbf{x}_{j}\right].$ (10)
It can be noted that there is a dependency between $\mathbb{E}_{q}[u]$ and $\mathbb{E}_{q}[\mathbf{h}_{j,a}]$, which is mitigated by using the previous iteration’s $\mathbb{E}_{q}[u]$ in the calculations, and in the case of it being the first iteration, it is randomly initiated.
@@ III-2 Maximization Step
This step involves determining the estimate of $\mathbf{x}_{j}$ by maximizing $\mathcal{F}(y_{j,a};q,\boldsymbol{\theta})$ with regards to $\boldsymbol{\theta}$, and it is shown in [zhang-deep-2022] that the transmitted signal from AP $j$, may be determined using
$\widehat{\mathbf{x}}_{j}=\operatorname*{argmin}_{\mathbf{x}_{j}}\left[(y_{j,a}-\mathbb{E}_{q}[\mathbf{h}_{j,a}]^{H}\mathbf{x}_{j})^{2}+M\mathbf{x}_{j}^{\,H}\boldsymbol{\Omega}_{\mathbf{h}}\mathbf{x}_{j}\right].$ (11)
@@ IV Privacy-Preserving Framework
For a cell-free MIMO ISAC system, we propose a joint, two block, precoder and AP selection framework for privacy-preservation, which utilizes alternating optimization of the two blocks. For a given receiver/transmitter AP configuration, the first block finds the precoding vectors that maximize the SINR at the sensing receivers. The second block minimizes the MI of the received signal at the users, and the sensing signal, by altering the AP configuration. Thereafter, a new set of precoders is generated and the AP configuration is optimized w.r.t. to the new precoder. As a convergence criteria, the procedure stops when two iterations leave the AP configuration unchanged. The first iteration assumes all APs as transmitters.
@@ IV-A Precoder Design
$P_{j,p}\leq P_{max},\forall j,\forall p,$ (12c)
where $\mathbf{W}$ is a $MN_{\text{Tx}}\times(N_{\text{UE}}+1)$ matrix consisting of each transmitter APs $M\times(N_{\text{UE}}+1)$ precoder matrix, denoted $\mathbf{W}_{j}$ for AP $j=1,2,...,N_{\text{Tx}}$ [demirhan_cell-free_2023]. In turn, $\mathbf{W}_{j}$ consists of AP $j$’s concatenated precoder vectors for each user plus the sensing precoder, that is $\mathbf{W}_{j}=\left[\mathbf{w}_{j,1},\mathbf{w}_{j,2},...,\mathbf{w}_{j,N_{\text{UE}}},\mathbf{w}_{j,t}\right]$. Hence, the aggregate precoder matrix $\mathbf{W}=\left[\mathbf{W}_{1}^{T},\mathbf{W}_{2}^{T},...,\mathbf{W}_{N_{\text{Tx}}}^{T}\right]^{T}$ columns represent the total precoding vector for a user from every transmitter AP, where the last column represents the target. Further, $\gamma_{\text{UE}_{i}}$ and $\gamma$ represents the total received SINR per user and the minimum SINR threshold for a user, respectively. $P_{j,p}$ is the transmit power from AP $j$’s antenna $p$, and $P_{max}$ is the maximum allowed transmission power for each antenna. Therefore, $P_{j,p}$ can be computed by taking the norm square of each row of $\mathbf{W}$.
@@
In (13), $N$ denotes the number of symbols that are being transmitted to each user, and, consequently, the number of samples in the sensing probing signal. Moreover, $\mathbf{s}[n]$ is the vector of transmitted symbols for all users and the sensing signal, i.e $\left[s_{1}[n],...,s_{N_{\text{UE}}}[n],s_{s}[n]\right]^{T}$.
Since the aim is to optimize $\gamma_{s}$ with respect to $\mathbf{W}$, the Concave-Convex Procedure (CCP) is considered and thus, the objective function must be concave [demirhan_cell-free_2023]. For this purpose the first order Taylor approximation is applied to the term $\mathbf{W}^{H}\mathbf{A}_{r}\mathbf{W}$ around the result of the previous iteration of the procedure, resulting in
@@
A similar approach is used to transform the communication SINR constraint into a convex function where the function $\gamma_{\text{UE}_{i}}$ becomes
@@
$\mathbf{w}_{i}=\left[\mathbf{w}^{T}_{1,i},\mathbf{w}^{T}_{2,i},...,\mathbf{w}^{T}_{N_{\text{Tx}},i}\right]^{T}$, representing the concatenated channel vector and precoder vector to user $i$ from all the transmitting APs, respectively. Again, the first order Taylor approximation is applied around the previous iteration’s value of $\mathbf{w}_{i}$. Also, the slack variables $\tau_{d}$ and $\tau_{n}$ for the denominator and numerator in (16), respectively, are introduced. The final expression for the constraint in (12b) is therefore